2.643 grams of potassium butanoate (KCH3(CH2)2CO2 ) is fully dissolved in 50.00 mL of
water, which is carefully transferred to a conical flask. Then 100.00 mL of 0.120 M HCℓ is
added dropwise to this solution from a burette.
1
Expert's answer
2020-09-04T10:16:11-0400

2.643 grams of potassium butanoate (KCH3(CH2)2CO2 ) is fully dissolved in 50.00 mL of

water .


Molar Mass of KCH3(CH2)2CO2  = 126 g / mol .


So , moles of KCH3(CH2)2CO2  = 2.643 / 126 = 0.02 mol ..


Given moles of HCl = 100 * 0.120 / 1000 = 0.0120 mol .


When HCl React with  potassium butanoate then   Butanoic Acid is formed .


So Hcl Is limiting Reagent( in lower amount ) So HCl Totally Consumed and  potassium butanoate and   Butanoic Acid Are Formed . this is buffer solution .


Moles of potassium butanoate =(0.02 - 0.0120) mol . = 0.008

Moles of butanoic acid = (0.0120) mol .


So , pH = pKa of butanoic acid + log [  potassium butanoate ] / [ Butanoic Acid ].


pH = 4.824 + log 0.008 / 0.0120 = 4.824 -0.174 = 4.65 Answer .